40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 10 15:41:37 2019, aaron, Update, IOO, WFS measurements Screenshot_from_2019-09-10_18-51-28.pngmctrend_190910.png
    Reply  Wed Sep 11 10:26:56 2019, aaron, Update, IOO, WFS measurements 
       Reply  Wed Sep 11 14:37:43 2019, aaron, Update, IOO, WFS measurements 20190911_WFS.jpg20190911_WFS_2.jpg
          Reply  Thu Sep 12 12:42:31 2019, aaron, Update, IOO, WFS measurements WFS2_TFs.pdfD755499D-9FDF-4E2B-BFC1-016B459DD35D.jpeg
             Reply  Fri Sep 13 10:36:03 2019, aaron, Update, IOO, WFS measurements TF-.pngWFS2_TFs.pdf
                Reply  Mon Sep 16 12:38:59 2019, aaron, Update, IOO, WFS measurements 
                Reply  Thu Oct 3 11:38:35 2019, aaron, Update, IOO, WFS measurements 
          Reply  Mon Sep 30 13:20:55 2019, aaron, Update, IOO, shot noise measurement 
       Reply  Fri Sep 13 10:53:40 2019, aaron, Update, IOO, WFS loop measurements 
          Reply  Mon Sep 16 05:08:04 2019, rana, Update, IOO, WFS loop measurements LSCsensingMatrix.ipynbASCsensingMatrix.ipynb
             Reply  Mon Sep 16 11:55:58 2019, rika, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 09:41:48 2019, gautam, Update, IOO, WFS loop measurements 
                Reply  Tue Sep 17 10:47:44 2019, rika, Update, IOO, WFS loop measurements MC2.pdf
                   Reply  Wed Sep 18 14:45:52 2019, rika, Update, IOO, WFS loop measurements Screenshot_from_2019-09-18_18-15-34.png
             Reply  Tue Sep 17 10:34:48 2019, aaron, Update, IOO, WFS loop measurements 
                Reply  Tue Oct 8 20:39:42 2019, aaron, Update, IOO, WFS loop measurements 
                   Reply  Wed Oct 9 09:37:28 2019, aaron, Update, IOO, WFS loop measurements 
Message ID: 14896     Entry time: Wed Sep 18 14:45:52 2019     In reply to: 14888
Author: rika 
Type: Update 
Category: IOO 
Subject: WFS loop measurements 

[aaron, rika]

Gettng TFs

In the data we got yesterday, we can see some filter's effect. 

But it is not good coherence above 10Hz, so we mesured again. And this time we save the data as xml file.

And also we chaned the frequency regions broader to watch corner frequency of suspension.

-----

 Diagnotics test tools

 range: 0.1 Hz to 100 Hz

 points: 120 

 Amplitude: 1000

----

but at low frequency, the mode maching cavity was unloked cause of too much shaking.

So, we saw single frequency TF, and searched the good amplitude.

 

First, I tried to get TF @0.1~1 Hz .

-----

0.1 to 1 Hz

points: 61 (I think it's too much becous it takes about an hour)

amplitude: 5

-----

The TFs and coherence of MC1/PIT to each QPD is below. [above window: coherence, below: TF]

During the mesurement, something happened @0.2-0.3Hz so I stopped it.

We found the coherence of WFS1P and WFS2Y is not good, but others are good.

we guess that it could come from alignment which made Q chainging to small.

 

Finaly, I also got the  .xml data of MC1P 1 Hz to 10 Hz. In this time,

-----

1 to 10 Hz

points: 41 

amplitude: 90

-----

 

Making matrics

Now we took single frequency 6 TFs (MC1/2/3 PIT/YAW) @7Hz (Because this frequency has good coherence in all channel).

Aaron wrote the script using dtt to making matrics. 

 

 

Quote:

[aaron, rika]

Once stop the auto-locker and realigned to make beam to get into QPD again.

After we lock MC, we took TFs from suspension MC1/2/3 PIT/YAW to WFS1/2 PIT/YAW. 

----- 

Diagnotics test tools

range: 7 Hz to 50 Hz

 avarage=61

Column 0: WFS2_PIT   1: WFS2_YAW   2:WFS1_PIT   3: WFS1_YAW   4: TRANCE_PIT   5:TRANCE_YAW 

-----

I'm wondering weather the MC1data I saved is correct, becouse I found the channel was changed when I exported MC2 data. So I took MC1 data again.

 

We got all data for TFs already.  Each data is devided to real part and imaginary part. Then we are arranging the datas to obtain TFs. 

TF of MC2 is attachiment 1. So tomorrow, I make other TF.

Quote:

[rika, aaron]

We aligned optics of WFS as it was. Now auto-locker is working to lock MC.

But it still doesn't lock. We notice that the c1lsc machine doesn't work. So we run rebootCILSC.sh.

 

Now we reset the hardware!

 

17:11

After reset, auto locking didn't work well. Gautum and Aaron reboot slow c1ioo. Then it works, and Gautam returned the MC to a good alignment.

We found the beam is not in the center of the QPD, we (turned off the MC autolocker and MC loop, then) realigned to make beam to get in to the QPD center. Afterwards we start auto locking.

With the WFS on, the maximum MC transmission we observe is 14,700 counts; after the transmission level stabilizes (MC_TRANS pit and yaw brought to 0), the MC transmission is only 14,200 counts. Perhaps the MC_TRANS QPD offsets need adjustment. We relieve the WFS servo of its DC offsets. This is the configuration we'll use for WFS loop measurements this week.

 

 

Attachment 1: Screenshot_from_2019-09-18_18-15-34.png  80 kB  Uploaded Wed Sep 18 19:27:41 2019  | Hide | Hide all
Screenshot_from_2019-09-18_18-15-34.png
ELOG V3.1.3-